OpenStudy (anonymous):

Choose the correct simplification of the expression (3x)4. 81x4 12x5 81x5 12x4

OpenStudy (jlastino):

Its equal to 3^4 x^4 or 3^2 3^2 x^4 you know what 3^2 is right?

OpenStudy (anonymous):

\[(xy)^a=x^ay^a\] what Jlastino said^

OpenStudy (anonymous):

yea 3^4 is 81. 3^2 is 9.

OpenStudy (anonymous):

where did you get 3^2 from ?

OpenStudy (anonymous):

from what jlastino said?

OpenStudy (anonymous):

wait so all i do is raise 3 to the forth power and x also the the forth power?

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

okay that helps, i had thought so but was unsure.

OpenStudy (anonymous):

so with (xyz^2)^4 i would do the same?

OpenStudy (anonymous):

yes, you just need to remember that \[\large(x^a)^b=x^{ab}\] for that z-term

OpenStudy (anonymous):

so i would come out with. x^4y^4z^16?

OpenStudy (anonymous):

no wait 8 for z

OpenStudy (anonymous):

right :) good job :)

OpenStudy (anonymous):

Awesome thanks :D

OpenStudy (anonymous):

very welcome :)

OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)2 I got the simplified expression of 9x^9y^16?